OpenStudy (anonymous):

45/4 (x^4) (y^3)

OpenStudy (anonymous):

Invert and multiply. Then you can cancel out terms as necessary.

OpenStudy (swag):

yes

OpenStudy (anonymous):

doesnt help...

OpenStudy (anonymous):

\[24y^5/15x^8 \div 8y^2/4x^4 = 24y^5/15x^8 \times 4x^4/8y^2\] \[24y^5/15x^8 \times 4x^4/8y^2 \] Then you can cancel and simplify.

OpenStudy (anonymous):

oops, copied and paste error. :P

OpenStudy (anonymous):

Ignore that 2nd line.

OpenStudy (anonymous):

so is it 4y^3/5x^4?